61. A 8000 Ω resistor is connected across 225 V. What current will flow? 62. What is wrong with Example 25-4 in the text? Explain.? 63.You need a typical cylinder resistor of 10% tolerance with a value of 13000 Ω. Draw a cylinder And indicate the first,second digits, multiplier and Tolerance colors on it?


please answer question 61 and show all the work
Answer #1

Resistance = R = 8000 ohm

Voltage = V = 225 volt

Current = I = V/R = 225/8000 = 0.028125 A = 28.125 mA